OpenStudy (anonymous):

even more sigma notations:

OpenStudy (anonymous):

\[\sum_{i=12}^{27}-5+13i\]

OpenStudy (anonymous):

@primeralph @waterineyes

OpenStudy (primeralph):

same thing from 12 to 27

OpenStudy (anonymous):

do i have to write it all out? is there an easier way or no?

OpenStudy (primeralph):

there are easier way.............but your professor might not want that

OpenStudy (primeralph):

*ways

OpenStudy (anonymous):

[-5+13(12)] etc. right?

OpenStudy (anonymous):

Just take the common things out of it..

OpenStudy (zarkon):

\[\sum_{a}^{b}[\alpha+\beta i]\] \[=\alpha(b-a+1)+\beta\left(\frac{b(b+1)}{2}-\frac{(a-1)a}{2}\right)\]

OpenStudy (anonymous):

[-5+13(12)]+[-5+13(13)]+[-5+13(14)]+[-5+13(15)]+[-5+13(16)]+ [-5+13(17)]+[-5+13(18)]+[-5+13(19)]+[-5+13(20)]+[-5+13(21)]+ [-5+13(22)]+[-5+13(23)]+ -5+13(24)]

OpenStudy (anonymous):

You will get like: 16*(-5) + 1[12 + 13 + 14 + 15........]

OpenStudy (primeralph):

well,,,,,,,,,,you get it

OpenStudy (anonymous):

That is 13 not 1 with those [] brackets..

OpenStudy (anonymous):

Otherwise, Zarkon has given you one formula, use it.. Even I write it on my notebook, I had not studied Sigma in my school much..

OpenStudy (anonymous):

[-5(x156]+ [-5(169)]+ [-5(182)]+ [-5(195)]+ [-5(208)]+ [-5(221)]+ [-5(234)]+ [-5(247)]+ [-5(260)]+ [-5(273)]+ [-5(286)]+ [-5(299)]+ [-5(312)]

OpenStudy (zarkon):

what calculator do you use?

OpenStudy (anonymous):

apple calculator

OpenStudy (zarkon):

ah...if you have a TI I could show you a way to do it on the calculator

OpenStudy (anonymous):

ok sry,lol

OpenStudy (zarkon):

you could also type sum[-5+13i,i=12..27] into wolfram alpha and get the answer

OpenStudy (anonymous):

-780 -845 -910 -975 -1040 -1105 -1170 -1235 -1300 -1365 -1430 -1495 -1560

OpenStudy (anonymous):

-15210

OpenStudy (anonymous):

that's not one of the choices

OpenStudy (anonymous):

i get it in wolfram alpha though